OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, -2) and has a slope of -3? x + 3y = 10 x + 3y = -10 3x + y = 10 3x + y = -10

OpenStudy (anonymous):

No lol. I can't remember anything that well since I just got out of surgery yesterday.

OpenStudy (anonymous):

(y+2)=-3(x-4)

OpenStudy (anonymous):

so then y=3(x-2)?

OpenStudy (anonymous):

so i subtract 3x to the other side and get y = -3x+10?

OpenStudy (anonymous):

Thank you so much